OpenStudy (boothepest1):

Find the two numbers that have a difference of 4 and a sum of 32. Numbers are____ and_____

OpenStudy (radar):

let x = one number let y = other number x - y = 4 given x + y = 32 given Solve.

jhonyy9 (jhonyy9):

just solve this system using elimination method do you know it ?

jhonyy9 (jhonyy9):

x-y=4 x+y=32 -------- 2x 0 =36 2x=36 x=36/2 x=? y=?